OpenStudy (anonymous):

integral of (ln x)^2

OpenStudy (alekos):

just go int[(lnx)^2]dx = int(2lnx)dx = 2int(lnx)dx

OpenStudy (alekos):

then you just work out the integral of lnx which is fairly straight forward

OpenStudy (anonymous):

Thanks!

OpenStudy (sidsiddhartha):

nope its not that easy

OpenStudy (anonymous):

But thats wrong...I think we need to integrate in parts

OpenStudy (sidsiddhartha):

yes we need by parts

OpenStudy (sidsiddhartha):

let \[u=[\ln(x)]^2\\and\\dv=dx ~~which ~~means\\ v=x\] now use by parts

OpenStudy (anonymous):

Thank you!

OpenStudy (sidsiddhartha):

no problem :) just use the formula and u'll get the answer

OpenStudy (alekos):

no. you don't need IBP

OpenStudy (alekos):

my apologies

OpenStudy (alekos):

you are correct. use IBP for the integral of lnx

OpenStudy (sidsiddhartha):

can u do it or need some help

OpenStudy (anonymous):

Got it. Thanks guys!

OpenStudy (gorv):

Integral ( [ln(x)]^2 dx ) use integration by parts. Let u = [ ln(x) ]^2. dv = dx. du = 2ln(x) (1/x) dx. Recall the formula of integration by parts: uv - Integral ( v du ) x [ln(x)]^2 - Integral ( x (2 ln(x)) (1/x) dx ) And we get some cancellations in the integral, particular with x and (1/x). x [ln(x)]^2 - Integral ( 2 ln(x) dx ) And let's factor out the 2 from the integral. x [ln(x)]^2 - 2 * Integral ( ln(x) dx ) From here, you would use integration by parts again. Let u = ln(x). dv = dx. du = (1/x) dx. v = x x [ln(x)]^2 - 2[ x ln(x) - Integral ( x(1/x) dx ) Distribute the -2, to get x [ln(x)]^2 - 2x ln(x) + 2 * Integral ( x(1/x) dx ) Simplify the integral, x [ln(x)]^2 - 2x ln(x) + 2 * Integral ( 1 dx ) And now it's an easy integral. With no more integrals to solve, don't forget to add a constant. x [ln(x)]^2 - 2x ln(x) + 2x + C And, to make the answer look like yours, let's factor the common monomial x from the three terms. x ( [ ln(x)]^2 - 2 ln(x) + 2 ) + C